# Curing Semantic Drift: A Dynamic Approach to Grounding Generation in Large Vision-Language Models

### v3

@@ 5.2 Experimental Results
$C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$
Nucleus 5.34 6.18 4.76 5.61 5.15 5.73
Our 7.73 6.87 7.38 6.75 6.64 6.56
VCD 5.55 6.38 5.32 6.09 5.55 5.97
Our 7.45 6.67 6.94 6.54 6.39 6.44
OPERA 6.16 6.59 6.02 6.17 5.98 5.93
Our 7.10 6.67 6.51 6.55 6.16 6.45
SID 5.69 6.54 5.34 6.07 5.55 6.05
Our 7.17 6.55 6.89 6.62 6.54 6.45
(b) Results on LLaVA-Bench-in-the-Wild [llava]
Settings LLaVA-1.5 InstructBLIP MiniGPT-4
$C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$
Nucleus 4.71 5.79 4.21 5.00 5.17 5.33
Our 8.42 7.04 7.25 6.42 6.00 5.75
VCD 5.25 5.96 4.88 5.54 5.00 5.08
Our 7.46 6.58 6.58 6.17 5.67 5.46
OPERA 5.75 6.75 5.88 5.62 5.75 6.08
Our 7.54 6.88 5.96 6.21 5.79 5.62
SID 5.58 6.79 4.42 5.25 5.38 5.42
Our 6.71 6.08 6.38 6.29 5.88 5.79
@@
Table 4: Robustness analysis of DLC with different CLIP backbones. Performance is evaluated on the CHAIR benchmark ($C_{S}\downarrow$ and $C_{I}\downarrow$) across three LVLMs.
@@
5.3 Ablation and Analysis
In this section, we conduct a series of experiments to systematically analyze our DLC framework. We validate the effectiveness of its core components, test its robustness to different visual backbones and its scalability to larger models. The Appendix provides additional experiments, including a hyperparameter sensitivity analysis (7.5) and results across different decoding strategies (7.8).
Robustness Analysis Across Different CLIP Models. To analyze the robustness of our DLC method and its dependence on the underlying CLIP model, we conducted comprehensive experiments using different CLIP backbone architectures. Specifically, we tested three models: the CLIP-336 (clip-vit-large-patch14-336) [clip], SigLIP (siglip-so400m-patch14-384) [siglip], and FG-CLIP (fg-clip-large) [fgclip]. The results in Table 2 and Table 4 demonstrate that DLC consistently improves performance across all combinations, validating its robustness.
Effectiveness of Core DLC Components. To isolate the contribution of each core component, we performed ablation studies evaluating model correctness and detailedness on the COCO [mscoco] dataset. The results are visualized in Figure 6. We validated the efficacy of: (1) the adaptive factor in the dynamic guidance strength $\lambda_{t}$, (2) the Current Contextual Token Alignment (CCTA), (3) the Isolated Token Alignment (ITA), and (4) the Relative Visual Advantage (RVA). As the figure clearly demonstrates, removing or replacing any of these components leads to a noticeable degradation in performance across all three tested LVLMs. These findings confirm that each component is integral to DLC’s effectiveness.
Methods
Max
Token
LLaVA-1.5-13B InstructBLIP-13B
$C_{S}$ $\downarrow$ $C_{I}$ $\downarrow$ $C_{S}$ $\downarrow$ $C_{I}$ $\downarrow$
Nucleus 64 23.2 8.16 27.8 10.2
DLC 64 18.2 5.50 24.2 7.75
Nucleus 512 53.2 15.1 64.0 19.6
DLC 512 33.0 8.80 55.6 15.4
Scalability to Larger Model Architectures. To validate DLC’s effectiveness on larger-scale models, we applied it to 13B-parameter versions of LLaVA-1.5 and InstructBLIP. The results in Table 5 show that DLC provides substantial improvements over the nucleus sampling baseline. This scalability is further confirmed in Appendix 7.7.
@@ 6 Conclusion
This paper confronted semantic drift—a dynamic failure where linguistic priors progressively hijack text generation from its visual grounding. We argued that existing training-free strategies are ill-equipped to solve this, being either computationally prohibitive, reliant on indirect proxies, or merely post-hoc corrections. To address this critical gap, we introduced DLC, the first framework to cure semantic drift in a direct, efficient, and fine-grained manner. Acting as a real-time visual referee, DLC employs a novel dual-aspect visual alignment check to compute a RVA against a dynamic baseline, which it then uses to surgically modulate the output logits. Extensive experiments demonstrated that DLC significantly reduces hallucinations and outperforms SOTA methods. Crucially, it achieves this robust performance while maintaining high inference efficiency, offering a practical and powerful path toward more reliable and trustworthy multimodal systems.
[…]
Supplementary Material
7 Supplementary Material
This appendix provides supplementary materials to support the main findings of our paper. We begin with additional background on LVLMs in Section 7.1. Section 7.2 provides further details on the benchmarks used in our experiments. We then present the detailed pseudo-code for DLC in Section 7.3 and elaborate on our specific implementation details in Section 7.4.
The remainder of the appendix is dedicated to extended experimental results and analyses. Section 7.5 presents a sensitivity analysis for our key hyperparameters, $\alpha$ and $N$. Section 7.6 provides the full results across all subtasks of the MME benchmark. Section 7.7 demonstrates DLC’s effectiveness on larger 13B-parameter models. Section 7.8 shows DLC’s compatibility with various other decoding strategies. Finally, Section 7.9 provides the exact prompts used for our GPT-4 and GPT-4o evaluations, and Section 7.10 presents additional qualitative case studies from the LLaVA-Bench [llava] dataset.
7.1 More backgrounds on LVLMs
The development of LVLMs has evolved from BERT-based [bert, vilbert, lxmert] cross-modal alignment to LLM-powered [flamingo, shikra, InstructBLIP, Blip-2, llava, minigpt4] architectures. Building upon foundational text-image alignment techniques like CLIP [clip] and BLIP [blip], contemporary LVLMs integrate visual encoders with LLM decoders through distinct optimization strategies. For instance, LLaVA [llava] employs lightweight adapters combined with instruction tuning, while InstructBLIP [InstructBLIP] and MiniGPT-4 [minigpt4] utilize Q-Former modules for visual token compression. Despite these innovations, significant challenges persist, particularly severe hallucination issues arising from misaligned multimodal representations. To systematically analyze mitigation strategies, we have examined three representative architectures, including LLaVA, InstructBLIP, and MiniGPT-4, whose results demonstrate that various technical approaches commonly share critical vulnerabilities to the hallucination issue.
7.2 Further Details on Benchmarks
To comprehensively evaluate our DLC method, we utilized several benchmarks specifically designed to assess hallucination in vision-language models. This section provides additional details about these benchmarks, including dataset characteristics, evaluation protocols, and metrics interpretation.
(1) CHAIR [chair] quantifies object hallucination in image captions by measuring the ratio of objects mentioned in captions but absent from ground-truth labels. It includes two variants: CHAIRI ($C_{I}$), calculating hallucinated objects versus all generated objects, and CHAIRS ($C_{S}$), assessing sentences with hallucinations versus total sentences. These two metrics are defined as follows:
$\begin{split}C_{I}&=\frac{|\{\text{hallucinated objects}\}|}{|\{\text{all mentioned objects}\}|},\\ C_{S}&=\frac{|\{\text{captions with hallucinated objects}\}|}{|\{\text{all captions}\}|}\end{split}$ (10)
Lower scores indicate better hallucination suppression.
(2) POPE [pope] (Polling-based Object Probing Evaluation) is a benchmark designed to directly and efficiently assess object-level hallucinations in LVLMs. It operates by querying the model with a series of simple polling-based (Yes/No) questions about the presence or absence of specific objects within an image. Based on the MS-COCO [mscoco] dataset, POPE generates both positive samples and, more importantly, challenging negative samples to rigorously test a model’s susceptibility to hallucination. Random: An object is randomly selected from all COCO categories to serve as a baseline test of factual grounding. Popular: An object with a high frequency of occurrence across the dataset is selected. This challenges models that rely heavily on statistical priors, tempting them to affirm the existence of common objects regardless of visual evidence. Adversarial: An object semantically, functionally, or spatially related to objects present in the image is chosen. For example, if the image contains a tennis racket, the model might be asked about a tennis ball. This probes for subtle, context-driven hallucinations where a model fabricates objects based on strong co-occurrence patterns. By evaluating the model’s accuracy, precision, recall, and F1-score across these question types, POPE provides a straightforward yet fine-grained metric for quantifying a model’s tendency to hallucinate.
Algorithm 1 Dynamic Logits Calibration (DLC)
1:Input image $v$, text prompt $x$, VLM $M_{\theta}$, CLIP model
2:Window size NN, modulation strength $\alpha$
3:Generated text sequence $y$
[…]
7:for $t=|x|+1$ to $|x|+\text{max\_new\_tokens}$ do
8:   $L_{t}\leftarrow M_{\theta}(y_{<t},v)$ $\triangleright$ Get original logits
9:   if $t>|x|+3$ then $\triangleright$ Start DLC after warm-up
10:     $y_{\text{ctx}}\leftarrow y_{t-N:t-1}$ $\triangleright$ Get preceding context
11:     $S_{t}^{\text{HCTA}}\leftarrow\text{CLIP}(v,y_{\text{ctx}})$ $\triangleright$ Calc. HCTA score
12:     Add $S_{t}^{\text{HCTA}}$ to $H$; maintain size NN
13:     $\bar{B}_{t}\leftarrow\operatorname*{avg}(H)$ $\triangleright$ Update historical baseline
14:     $C\leftarrow\operatorname*{TopK}(L_{t})$ $\triangleright$ Get top-k candidate tokens
15:     for each candidate token $c_{i}\in C$ do
16:        $S_{t,i}^{\text{CCTA}}\leftarrow\text{CLIP}(v,y_{\text{ctx}}\oplus c_{i})$
17:        $S_{t,i}^{\text{ITA}}\leftarrow\text{CLIP}(v,c_{i})$
18:        $S_{t,i}^{\text{COMB}}\leftarrow(S_{t,i}^{\text{CCTA}}+S_{t,i}^{\text{ITA}})/2$
19:        $\text{RVA}_{t,i}\leftarrow(S_{t,i}^{\text{COMB}}-\bar{B}_{t})/(1-\bar{B}_{t})$
20:        $\lambda_{t}\leftarrow\alpha\cdot(1-\bar{B}_{t})^{2}$ $\triangleright$ Dynamic strength
21:        $L^{\prime}_{t,i}\leftarrow L_{t,i}\cdot\exp(\lambda_{t}\cdot\sigma(\text{RVA}_{t,i}))$ $\triangleright$ Adjust logit
22:     end for
23:     $L_{t}\leftarrow L^{\prime}_{t}$ $\triangleright$ Update logits with adjusted ones
24:   end if
25:   $y_{t}\leftarrow\text{sample}(L_{t})$
26:   $y\leftarrow y\oplus y_{t}$
27:   if $y_{t}$ is end token then break
28:   end if
29:end for
30:return $y$
(3) SHR Evaluations [hadpo] function as a crucial complement to conventional assessment methodologies. While traditional metrics such as CHAIR effectively quantify hallucinations at the object-existence level, they demonstrate inadequacy in detecting more fine-grained hallucination types, including positional, relational, and attribute-based inaccuracies. To address this limitation, GPT-4 assisted evaluation employs detailed object-level descriptions from the Visual Genome (VG) [vgdataset] dataset as ground-truth reference, leveraging GPT-4’s sophisticated capabilities for precise hallucination detection and classification. The complete prompt template is provided in the Appendix 7.9.
(4) GPT-4o assisted evaluations [^1] We utilize GPT-4o for assisted evaluation, as it offers enhanced capabilities over previous vision-capable models like GPT-4V, which is also being phased out by OpenAI. [woodpecker] further conduct a comprehensive analysis of hallucinations and text quality in open-ended generation tasks. Harnessing GPT-4o’s sophisticated capabilities, [woodpecker] established a numerical assessment protocol utilizing a 0-10 scale, which evaluates two essential dimensions: Accuracy-the extent to which responses align with image content without fabrication, and Detailedness-the degree of informational richness present in the responses.
(5) MME Benchmark [mme] provides a comprehensive assessment of LVLM capabilities by evaluating both perceptual and cognitive skills. The benchmark is structured into a total of 14 subtasks, with ten designed to test perception and four to assess cognition. All evaluations are performed using a yes/no question format. MME features specific sub-benchmarks for hallucination discrimination, including tasks that probe existence, count, position, and color. These tasks are intended to examine both object-level and attribute-level hallucinations. Task scores are used as the metric to report performance on this benchmark.
(6) LLaVA-Bench-in-the-Wild [llava] is designed to evaluate LVLMs’ proficiency in addressing complex tasks and their adaptability to novel domains. We perform targeted case studies on this dataset to demonstrate the efficacy of our proposed DLC.
7.3 Pseudo-code of Our Method
In this section, we provide a detailed pseudo-code of our Dynamic Logits Calibration (DLC) method, as shown in Algorithm 1.
7.4 Implementation Details
In this section, we provide comprehensive implementation details of our Dynamic Logits Calibration (DLC) method, including model configuration, parameter settings, and computational workflows. These details are essential for reproducing our results and understanding the practical aspects of DLC.
Unless otherwise specified, we compute all visual alignment scores using the SigLIP model [siglip]. The historical baseline $\bar{B}_{t}$ is calculated over a sliding window of $N=8$ past context scores, following a 3-step warm-up period. For the top-k candidates ($k=50$) at each step, the combined score $S_{t,i}^{\text{COMB}}$ is computed. The adaptive modulation employs a maximum intervention strength $\alpha=3$, and the standard sigmoid function $\sigma(\cdot)$ for logit adjustment.
CLIP Query Efficiency. To ensure high throughput, the $k=50$ CLIP queries required at each decoding step are implemented for high efficiency. This efficiency stems from two key factors: (1) Batch Processing: All 50 candidate queries are batched into a single, parallel forward pass through the CLIP text encoder, which is highly optimized on modern GPUs. (2) Low Computational Load: The text input for each query is extremely short. The computational cost of encoding these short snippets is negligible compared to the cost of a full forward pass through the multi-billion parameter LVLM. As a result, this process has a minimal impact on inference speed. This single, lightweight CLIP pass is a core design choice for efficiency, contrasting with methods like VCD [vcd] or SID [sid] that require multiple, costly forward passes through the entire LVLM.
7.5 Sensitivity Analysis of Key Hyperparameters: $\alpha$ and NN
The effectiveness of our DLC method hinges on two key hyperparameters: the modulation strength $\alpha$ and the context window size $N$. This section presents a sensitivity analysis of these parameters, evaluating their impact on hallucination mitigation across three distinct LVLMs. Performance is measured using the CHAIR metric [chair], with a maximum of 512 new tokens generated for all ablation studies.
Impact of Modulation Strength ($\alpha$): We first examine the modulation strength, $\alpha$. Our findings, detailed in Table 6, reveal that the optimal $\alpha$ setting is highly dependent on both the LVLM and the decoding strategy.
With Greedy Search, lower $\alpha$ values generally yield better performance. Specifically, $\alpha=1$ delivers the strongest results for LLaVA-1.5 and InstructBLIP. For MiniGPT-4, performance is competitive at both low ($\alpha=1$) and moderate ($\alpha=3$) strengths.
With Nucleus Sampling, the trend reverses, with moderate-to-high $\alpha$ values (from 3 to 5) consistently improving CHAIR scores. Optimal performance is observed around $\alpha=4$ for LLaVA-1.5, $\alpha=3$ for MiniGPT-4, and $\alpha=5$ for InstructBLIP.
These results underscore that while a global value of $\alpha=3$ is used for our main experiments, tailoring this parameter to the specific model and decoding method can unlock further performance gains.
Impact of Context Window Size ($N$): Next, we analyze the influence of the context window size, $N$. As shown in Table 7, the optimal choice of $N$ also varies significantly across different configurations.
With Greedy Decoding, InstructBLIP and MiniGPT-4 generally benefit from larger window sizes, achieving their best scores at $N=16$. In contrast, the effect on LLaVA-1.5 is less monotonic, with smaller to moderate window sizes ($N=4$ or $N=12$) proving more advantageous.
With Nucleus Sampling, the response is more varied. MiniGPT-4 performs best with a small window ($N=4$), showing a decline in performance as $N$ increases. LLaVA-1.5’s performance fluctuates without a clear trend, while InstructBLIP shows no significant improvement from changes in $N$.
While we adopt a uniform $N=8$ for our main experiments, this analysis highlights that model- and strategy-specific tuning of the window size is crucial for maximizing performance.
In summary, both modulation strength $\alpha$ and window size $N$ are influential hyperparameters for the DLC method. The ablation studies presented here clearly demonstrate that while the default values used in this paper provide a strong general baseline, optimal results are achieved by tuning these parameters for each LVLM and decoding strategy combination.
Table 6: Impact of Modulation Strengths $\alpha$ on CHAIR Metrics for Different LVLMs and Decoding Strategies
Method $\alpha$ LLaVA-1.5 InstructBLIP MiniGPT-4
Greedy 1 33.8 9.85 47.9 13.2 32.6 10.3
2 34.1 10.0 48.1 13.5 33.7 10.1
3 34.8 10.1 49.6 13.5 33.0 9.89
4 35.3 9.87 50.5 14.2 33.5 10.0
5 36.8 10.5 51.0 14.4 33.9 10.6
Nucleus 1 41.8 12.7 55.4 15.4 34.6 11.1
2 40.3 11.4 53.2 14.8 34.0 10.6
3 39.4 10.9 51.0 14.3 32.0 11.2
4 37.5 10.8 50.8 14.7 33.6 10.7
5 37.5 10.9 50.6 14.5 33.6 10.9
Table 7: Impact of Window Size NN on CHAIR Metrics for Different LVLMs and Decoding Strategies
Method NN LLaVA-1.5 InstructBLIP MiniGPT4
Greedy 4 34.4 9.45 52.2 14.6 35.4 11.1
8 35.1 9.72 49.0 12.8 33.4 10.3
12 34.4 9.51 48.0 13.5 33.2 9.85
16 37.0 9.84 47.0 13.2 31.0 9.69
Nucleus 4 38.6 11.1 51.4 14.3 31.0 10.2
8 39.4 10.9 51.0 14.3 32.0 11.2
12 38.6 11.5 54.8 14.8 32.4 10.1
16 37.6 11.4 50.4 15.0 34.0 12.6
7.6 Full MME Benchmark Results
As discussed in the main paper, the MME benchmark [mme] offers a comprehensive evaluation of an LVLM’s capabilities across 14 subtasks spanning perception and cognition. While Figure 5 in the main text highlights key findings regarding hallucination-related subtasks, this section provides the complete performance breakdown of DLC compared to baseline methods across all MME subtasks. The detailed scores, presented in Figure 7, further substantiate our claim that DLC effectively mitigates hallucinations without degrading the model’s foundational multimodal abilities.
Figure 7: Evaluation on the comprehensive MME benchmark. The results on LLaVA-1.5 7B demonstrate that DLC preserves the model’s general-purpose capabilities across the benchmark’s 14 subtasks spanning perception and cognition.
7.7 Scalability to Larger Model Architectures
To complement the CHAIR evaluation, we also conducted GPT-4o assisted evaluation focusing on correctness and detailedness of generated descriptions. The results are presented in Table 8. The GPT-4o evaluation demonstrates substantial improvements in both correctness and detailedness.
Table 8: GPT-4o Evaluation Results on 13B Models.
Methods LLaVA-1.5-13B InstructBLIP-13B
$C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$
Nucleus 5.25 6.05 4.26 5.06
DLC 7.75 6.77 7.00 6.36
7.8 DLC Performance with Various Decoding Strategies
Our proposed DLC demonstrates compatibility with various mainstream decoding strategies. To comprehensively evaluate its effectiveness, in addition to its performance with common strategies such as nucleus sampling and greedy sampling, we conducted further experiments on the LLaVA-1.5 [llava] model using the MSCOCO dataset [mscoco] and the CHAIR benchmark [chair]. These experiments explored other decoding methods and parameters, specifically: Top-p sampling (p=0.9), Top-k sampling (k=50), and Top-k sampling (k=50) with varying temperatures (t=1.5 and t=0.8). The results, presented in Figure 8, consistently indicate that regardless of the sampling method employed, DLC effectively reduces hallucinations and enhances overall model performance. This substantiates the robustness and effectiveness of our proposed DLC across diverse decoding strategies.
Figure 8: Results of different decoding strategies.
7.9 Prompts for GPT-4 and GPT-4o Evaluation
In this section, we provide the prompts for GPT-4 and GPT-4o evaluation, as shown in Figure 9 and Figure 10.
Figure 9: Prompts of GPT-4 for evaluations.
Figure 10: Prompts of GPT-4o for evaluations.
7.10 Extended Comparative Results on LLaVA-Bench
In this section, we present additional comparative results of our DLC approach against other methods on the LLaVA-Bench [llava]. These extended evaluations provide a more comprehensive understanding of our method’s performance across various question types and scenarios.
Figure 11: Case Study: LLaVA-1.5 and InstructBLIP on LLaVA-Bench-in-the-Wild (Hallucinations Marked Red).
Figure 12: Case Study: LLaVA-1.5 and InstructBLIP on LLaVA-Bench-in-the-Wild (Hallucinations Marked Red).

### v4

@@ 5.2 Experimental Results
COCO L-Bench COCO L-Bench COCO L-Bench
$C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$
Nucleus 5.34 6.18 4.71 5.79 4.76 5.61 4.21 5.00 5.15 5.73 5.17 5.33
Our 7.73 6.87 8.42 7.04 7.38 6.75 7.25 6.42 6.64 6.56 6.00 5.75
VCD 5.55 6.38 5.25 5.96 5.32 6.09 4.88 5.54 5.55 5.97 5.00 5.08
Our 7.45 6.67 7.46 6.58 6.94 6.54 6.58 6.17 6.39 6.44 5.67 5.46
OPERA 6.16 6.59 5.75 6.75 6.02 6.17 5.88 5.62 5.98 5.93 5.75 6.08
Our 7.10 6.67 7.54 6.88 6.51 6.55 5.96 6.21 6.16 6.45 5.79 5.62
SID 5.69 6.54 5.58 6.79 5.34 6.07 4.42 5.25 5.55 6.05 5.38 5.42
Our 7.17 6.55 6.71 6.08 6.89 6.62 6.38 6.29 6.54 6.45 5.88 5.79
Broad Capabilities on MME. To verify that our hallucination mitigation strategy does not compromise general multimodal capabilities, we conducted a comprehensive evaluation on the MME benchmark [mme]. The results, illustrated in Figure 4, reveal a nuanced trade-off. In Perception tasks, DLC achieves consistent gains, corroborating its effectiveness in correcting visual grounding errors. In Cognition tasks, performance varies by type. For visually-grounded reasoning, DLC improves over the baseline, as clearer perception aids understanding. However, for abstract symbolic tasks such as Code Reasoning and Numerical Calculation, we observe a minor performance regression compared to the unconstrained base model. This suggests that enforcing strict visual alignment may impose a slight constraint on pure symbolic reasoning. Notably, however, DLC avoids the catastrophic degradation seen in contrastive baselines, maintaining a much more balanced capability profile.
Figure 4: Comprehensive breakdown of MME benchmark performance on LLaVA-1.5. We compare DLC variants against standard decoding baselines across 14 subtasks spanning perception and cognition suites.
5.3 Ablation and Analysis
Table 4: Efficiency benchmarking and sensitivity analysis. We report end-to-end inference time (s/sample) and GPU memory usage on LLaVA-1.5 (NVIDIA RTX 3090). The upper section compares DLC with standard baselines; the lower section analyzes the trade-off between candidate pool size $k$, latency, and faithfulness.
Method Time $\downarrow$ Memory (MB) $\downarrow$ CHAIRs $\downarrow$ CHAIRi $\downarrow$
Vanilla 3.64 15062 56.6 18.2
VCD 5.22 15062 59.4 18.9
OPERA 7.43 20854 53.7 15.0
DLC (k=10k{=}10) 4.84 17874 41.2 11.2
DLC (k=30k{=}30) 5.11 17874 38.2 10.6
DLC (k=50k{=}50) 5.35 17868 35.8 10.5
DLC (k=100k{=}100) 6.25 17862 36.6 10.3
DLC (k=200k{=}200) 9.09 18118 39.8 11.5
In this section, we present a systematic analysis of DLC. Specifically, we (i) ablate key components to verify their individual contributions, (ii) study the effect of the candidate pool size $k$ and its quality–cost trade-off in runtime and memory, and (iii) evaluate robustness to different visual backbones and scalability to larger LVLMs. We include further results in the Appendix, such as hyperparameter sensitivity and evaluations under different decoding strategies.
Effectiveness of Core Components. To isolate the contribution of each core component, we performed ablation studies evaluating model correctness and detailedness on the COCO dataset. The results are visualized in Figure 5. We validated the efficacy of: (1) the adaptive factor in the dynamic guidance strength $\lambda_{t}$, (2) the Current Contextual Token Alignment (CCTA), (3) the Isolated Token Alignment (ITA), and (4) the Relative Visual Advantage (RVA). As the figure clearly demonstrates, removing or replacing any of these components leads to a noticeable degradation in performance across all three tested LVLMs. These findings confirm that each component is integral to DLC’s effectiveness.
Efficiency and Sensitivity Analysis. Table 4 presents a dual analysis: comparing DLC’s efficiency against established baselines and dissecting the impact of the candidate pool size $k$. DLC demonstrates a superior efficiency-quality profile. Notably, at a moderate setting, DLC is faster than VCD while drastically reducing hallucination. Meanwhile, the ablation reveals a decoupled scaling behavior. Inference time increases linearly with $k$ due to the computational cost of scoring more candidates. However, GPU memory usage remains remarkably stable regardless of $k$. This confirms that the memory overhead is static rather than dynamic, making DLC memory-friendly even with larger candidate pools. Furthermore, performance does not scale monotonically. While increasing $k$ from 10 to 50 significantly improves grounding by capturing more relevant candidates, pushing $k$ to extremes yields diminishing returns and even degrades performance. This degradation likely stems from long-tail noise—introducing too many low-probability tokens that confuse the selection mechanism. Consequently, we recommend $k\in[30,50]$ as the optimal range, offering state-of-the-art faithfulness with inference speeds comparable to or faster than contrastive baselines.
Methods Max Token LLaVA-1.5 InstructBLIP
$C_{S}$ $\downarrow$ $C_{I}$ $\downarrow$ $C_{S}$ $\downarrow$ $C_{I}$ $\downarrow$
Nucleus 64 23.2 8.16 27.8 10.2
DLC 64 18.2 5.50 24.2 7.75
Nucleus 512 53.2 15.1 64.0 19.6
DLC 512 33.0 8.80 55.6 15.4
Table 6: GPT-4o Evaluation on 13B Models. Assessment of correctness and detailedness.
Methods LLaVA-1.5 InstructBLIP
$C\uparrow$ $D\uparrow$ $C\uparrow$ $D\uparrow$
Nucleus 5.25 6.05 4.26 5.06
DLC 7.75 6.77 7.00 6.36
VCD 5.93 6.44 5.27 5.80
DLC 7.32 6.63 6.86 6.09
@@ 5.3 Ablation and Analysis
Table 7: Robustness analysis of DLC with different CLIP backbones. Performance is evaluated on the CHAIR benchmark across three LVLMs.
@@
Robustness Analysis Across Different CLIP Models. To analyze the robustness of our DLC method and its dependence on the underlying CLIP model, we conducted comprehensive experiments using different CLIP backbone. Specifically, we tested three models: the CLIP-336 (clip-vit-large-patch14-336) [clip], SigLIP (siglip-so400m-patch14-384) [siglip], and FG-CLIP (fg-clip-large) [fgclip]. The results in Table 2 Table 7 and Figure 4 demonstrate that DLC consistently improves performance across all combinations, validating its robustness.
Scalability to Larger Model Architectures. We further evaluate DLC on 13B variants of LLaVA-1.5 and InstructBLIP (Tables 6 and 6). DLC consistently outperforms baselines across both metrics: it significantly reduces object hallucination in CHAIR evaluation while simultaneously improving holistic correctness and detailedness scores in GPT-4o assessments. These findings confirm that DLC generalizes effectively to stronger backbones without specific tuning.
@@ 6 Conclusion
This paper tackles semantic drift—a dynamic failure mode where linguistic priors progressively hijack LVLM generation and detach it from visual evidence. We argue that hallucination is best understood as a token-level trajectory failure: as drift emerges, the model can still place grounded candidates among the top options, yet selects a linguistically plausible but visually unfaithful token. To address this root cause, we introduced DLC, a training-free decoding framework that injects explicit visual grounding signals into token selection in real time. DLC acts as a visual referee by combining intrinsic relevance and contextual coherence checks and comparing them to an adaptive historical baseline, thereby computing a relative visual advantage used to dynamically calibrate logits. Across multiple LVLMs and diverse benchmarks, DLC consistently reduces hallucinations while maintaining response quality and broad multimodal capabilities. Further analyses on different vision backbones and the candidate pool size reveal a favorable quality–cost trade-off with predictable overhead. We hope our findings on semantic drift and our decoding-time grounding mechanism will inspire future work on more reliable and trustworthy multimodal generation.